# Laboratoire de Physique Théorique et Hautes Énergies (LPTHE) Sorbonne Université & CNRS, UMR 7589 F-75005 Paris, France

### v1

@@
We introduce $\mathrm{U}(N)$ invariant polynomials on the space of $N\times N$ matrices that are precursors of free cumulants in various respects. First, they are polynomials of deterministic matrices, that are not yet evaluated over some probability law, contrary to what is usually meant by cumulants. Secondly, they converge towards the algebraic expression of free cumulants in terms of moments as $N\to\infty$, with $1/N^{2}$ corrections expressed in terms of monotone Hurwitz numbers. Their most crucial property is their additivity with respect to averaging over sums of $\mathrm{U}(N)$ conjugacy orbits, providing a finite $N$ version of the well-known additivity of free cumulants in free probability. Finally, they extend several properties of free cumulants at finite $N$, including a Wick rule for their average over a Gaussian weight and their appearance in various matrix integrals. Building on the additivity property of these precursors, we also define and compute a coproduct describing the behaviour of general invariant polynomials with respect to the addition of $\mathrm{U}(N)$ conjugacy orbits, as well as their expectation values on sums of $\mathrm{U}(N)$-invariant random matrices. In our construction, a central role is played by the so-called HCIZ integral, both for the definition of the precursors and for the derivation of their properties.
@@ 1.1 Free cumulants and their additivity
where $DU_{N}$ stands for the Haar measure on $\mathrm{U}(N)$. The question then naturally arises: is there a finite $N$ analogue of the free cumulants $\kappa_{n}$ that satisfies this additivity property? In this article, we will define such “precursors” of the free cumulants and will study their properties in detail, as summarised in the next section. We also refer to section 1.3 for a brief review and discussion of pre-existing works in this direction.
@@ 1.3 A short review of related work
  • In a recent work[14], mostly devoted to random tensors, a concept of cumulants for finite tensors is developed. When applied to 2-tensors regarded as matrices, these objects seem to coincide with ours. Our construction, however, proceeds along a very different route (the HCIZ integral), which enables us to derive the many results of Theorem 1. It would be interesting to study the relation between these two approaches in more details.
@@ 2.2 Generalised precursors of free cumulants
The above expressions for $K_{\alpha}(A)$ and $K_{n}(A)$ form the point 2 of our main Theorem 1.
@@ 3.2 $\boldsymbol{1/N^{2}}$ corrections and Hurwitz numbers
The prefactors in equations (69) and (70) might seem slightly strange at first sight, as they break the symmetry $\alpha\leftrightarrow\beta$. However, they have been choosen to ensure a remarkable property (see for instance [30, Lemma 2.9]), namely the fact that the matrices $\bigl{(}H^{\bullet\hskip 1.0pt,\hskip 1.0pt<}(\alpha,\beta)\bigr{)}_{\alpha,\beta\vdash n}$ and $\bigl{(}H^{\bullet\hskip 1.0pt,\hskip 1.0pt\leq}(\alpha,\beta)\bigr{)}_{\alpha,\beta\vdash n}$ are inverse one of another, i.e.,
@@ 4.4 Relation between cyclic products and generalised precursors
for any $n$-plet of pairwise distinct indices $i_{1},i_{2},\cdots i_{n}$ and no summation over repeated indices. This relation seems to have been pointed out first in [6, Theorem 2.6], see also [7, Theorems 1 and 2]. An elegant proof has been given in a recent paper of Bernard and Hruza [8]. In fact their proof extends word for word to finite $N$, substituting the precursor $K_{n}$ for the free cumulant $\kappa_{n}$:
@@ 5.3 An illustration: Horn’s problem
Figure 1: A numerical experiment with a sample of 200,000 $4\times 4$ matrices $UBU^{\dagger}, AA$ and $B$ of regularly spaced spectrum on [1,1][-1,1], showing the histograms of $\delta\kappa_{4}(U)$ (left), $\delta K_{4}(U)$ (middle) and $\delta f_{4}(U)$ (right). The improvement of the latter two with respect to the former is manifest: smaller skewness, distribution closer to Gaussian, etc. The broader range of values of $\delta K_{4}$ with respect to $\delta\kappa_{4}$ is due to the factor $N^{4}(N^{2}+1)/((N^{2}-1)(N^{2}-4)(N^{2}-9)$ in (55).
@@ 5.4 Behaviour of more general polynomials and the orbit coproduct
The algebra $\mathcal{A}^{(N)}$ can also be seen as the algebra of symmetric polynomials in $N$ variables, by passing to eigenvalues of matrices. We note that the algebra of symmetric polynomials in an arbitrary number of variables is equipped with a well-known cocommutative coassociative graded coproduct, defined by splitting any such polynomial into a sum of products of polynomials in two subsets of the variables (see for instance [35, Example I.5.25]). Although this coproduct shares some common properties with the one $\Delta$ studied here, it also differs from it on various aspects. To start with, $\Delta$ is defined on symmetric functions of a fixed number $N$ of variables and will depend explicitely on $N$, in contrast with this other coproduct. Moreover, the latter defines a Hopf algebra structure whereas we observed above that $\Delta$ does not. We will comment on other similarities and differences of these two coproducts throughout this subsection.
@@ 6.3 Collins–Gurau–Lionni cumulants and additive functionals
where $A$ is a random matrix which is integrated along a given measure $\mu\in\mathcal{M}^{(N)}$, while $B$ is kept deterministic and will serve as a bookkeeping variable. The Collins-Gurau-Lionni (CGL) cumulants $K^{\text{CGL}}_{\alpha}[\mu]$ are then extracted from the expansion of the logarithm of this quantity along Newton polynomials of $B$ – see [15, Equation (4.22)]: [^14] Here, we have expanded $\log\overline{Z}_{B,z}[\mu]$ along the Netwon basis $\{p_{\alpha}(B)\}_{\alpha\in\mathcal{P}}$, as defined in Section 5.4 (see above equation (113)). We have used a different normalisation than [15], which simplifies some of the discussions below.
@@
The main property of the CGL cumulants is their additivity with respect to the convolution $\star$, as established in [15, Proposition 4.9]. Namely, we have
@@ Discussion.
More generally, the normalisation in equation (144) has been chosen such that the CGL cumulants $K^{\text{CGL}}_{\alpha}$ have a finite and non-zero $N\to\infty$ limit. More precisely, according to [15, Equation (3.27)], $\lim_{N\to\infty}K^{\text{CGL}}_{\alpha}[\mu]$ becomes the higher-order free cumulant [6] associated with the partition $\alpha=(\alpha_{1},\dots,\alpha_{\ell})$ (in a certain normalisation). This is consistent with the addivity property (149) of $K^{\text{CGL}}_{\alpha}$, since higher-order free cumulants are known to be additive quantities with respect to the free convolution.
@@
$\mathbb{E}_{A\sim\mu}(A_{11}^{n})=\sum_{\alpha\vdash n}|\text{Cl}_{\alpha}|\,N^{\ell(\alpha)-n}\overline{K}_{\alpha}[\mu]\,$, (157)
@@ 7.1 Finite $\boldsymbol{N}$ free convolution associated with precursors?
$P_{A}(x)=\det(x\,\mathbb{I}-A)=\prod_{i=1}^{N}(x-\alpha_{i})=\sum_{k=0}^{N}(-1)^{k}x^{n-k}e_{k}(A)\,.$ (161)
@@ Acknowledgments
It is a pleasure to acknowledge stimulating discussions with D. Bernard, J. Bouttier, É. Brézin, P. Di Francesco, M. Maïda, S. Tarricone and P. Zinn-Justin. We are particularly indebted to P. Biane for drawing our attention to the point discussed in Section 7.1.
@@ Appendix D Strictly monotone Hurwitz numbers and counting permutations by genus
We will need some notations and preliminary results. For $r\in\{1,\dots,n-1\}$, we let $\Sigma_{n,r}$ be the set of $r$-uples $(\tau_{1},\dots,\tau_{r})$ of transpositions in $S_{n}$ such that $\tau_{i}=(a_{i}\,b_{i})$ with $a_{i}<b_{i}$ and $b_{1}<\cdots<b_{r}$. If a permutation $\sigma\in S_{n}$ can be written as a product $\sigma=\tau_{1}\cdots\tau_{r}$ with $(\tau_{1},\dots,\tau_{r})\in\Sigma_{n,r}$, then we say that $\tau_{1}\cdots\tau_{r}$ is a primitive factorisation of $\sigma$ of length $r$. The following result is standard (see for instance [45, Lemma 1.5.1]) and will play a crucial role in our proof.

### v2

@@
We study $\mathrm{U}(N)$ invariant polynomials on the space of $N\times N$ matrices first introduced by Capitaine and Casalis, that are precursors of free cumulants in various respects. First, they are polynomials of deterministic matrices, that are not yet evaluated over some probability law, contrary to what is usually meant by cumulants. Secondly, they converge towards the algebraic expression of free cumulants in terms of moments as $N\to\infty$, with $1/N^{2}$ corrections expressed in terms of monotone Hurwitz numbers. Their most crucial property is their additivity with respect to averaging over sums of $\mathrm{U}(N)$ conjugacy orbits, providing a finite $N$ version of the well-known additivity of free cumulants in free probability. Finally, they extend several properties of free cumulants at finite $N$, including a Wick rule for their average over a Gaussian weight and their appearance in various matrix integrals. Building on the additivity property of these precursors, we also define and compute a coproduct describing the behaviour of general invariant polynomials with respect to the addition of $\mathrm{U}(N)$ conjugacy orbits, as well as their expectation values on sums of $\mathrm{U}(N)$-invariant random matrices. In our construction, a central role is played by the so-called HCIZ integral, both for the construction of the precursors and for the derivation of their properties.
@@ 1.1 Free cumulants and their additivity
where $DU_{N}$ stands for the Haar measure on $\mathrm{U}(N)$. The question then naturally arises: is there a finite $N$ analogue of the free cumulants $\kappa_{n}$ that satisfies this additivity property? Such “precursors” of the free cumulants were defined by Capitaine and Casalis in [6]. In this article, we will propose a new approach to construct and study these precursors based on the HCIZ integral, allowing us to recover various results on their properties and derive new ones, summarised in the next section. We also refer to section 1.3 for a brief review and discussion of pre-existing works in related directions.
@@ 1.3 A short review of related work
  • The precursors of free cumulants $K_{n}$ and their generalisations $K_{\alpha}$ (or rather their expectation values $\overline{K}_{\alpha}$) were first introduced by Capitaine and Casalis in [6]. In that reference, they are defined using the convolution on the symmetric group (see also [10] for a more geometric interpretation). In contrast, our construction is based on the HCIZ integral. We recover the expression of $K_{\alpha}$ as a convolution in equation (47), thus making the link with [6] explicit. Some of the properties of the generalised precursors summarised in our main Theorem 1 were already found in [6, 10], including the additivity of $K_{n}$ (point 3), their large $N$ limit (point 4) and their expectation value under a Gaussian distribution (point 8). Our approach allows us to rederive these results from a different route and to obtain various new ones, in particular the relation between the large-$N$ expansion of $K_{\alpha}$, Hurwitz numbers and the enumeration of permutations by genus (points 5 and 6 in Theorem 1), as well as the generating function of $K_{n}$ (point 7) and the behaviour of general invariant polynomials with respect to sums of conjugacy orbits (see Sections 5.4 and 6).
@@
  • In a recent work[16], mostly devoted to random tensors, a concept of cumulants for finite tensors is developed. When applied to 2-tensors regarded as matrices, these objects seem to coincide with the generalised precursors $K_{\alpha}$. It would be interesting to study the relation between their construction and that of [6, 10] or the present paper.
@@ 2.2 Generalised precursors of free cumulants
The above expressions for $K_{\alpha}(A)$ and $K_{n}(A)$ form the point 2 of our main Theorem 1. The right-hand side of (47) can be rephrased as the convolution over the symmetric group $S_{n}$ of the Weingarten coefficient $\sigma\in S_{n}\mapsto\mathrm{Wg}([\sigma])$ with the function $\sigma\in S_{n}\mapsto p_{[\sigma]}(A)$. Up to a global power of $N$, we thus recover the matrix cumulant introduced by Capitaine and Casalis in [6, 10], showing that our construction of the generalised precursor from the HCIZ integral coincides with their definition as a convolution over $S_{n}$.
@@ 3.2 $\boldsymbol{1/N^{2}}$ corrections and Hurwitz numbers
The prefactors in equations (69) and (70) might seem slightly strange at first sight, as they break the symmetry $\alpha\leftrightarrow\beta$. However, they have been choosen to ensure a remarkable property (see for instance [32, Lemma 2.9]), namely the fact that the matrices $\bigl(H^{\bullet\hskip 1.0pt,\hskip 1.0pt<}(\alpha,\beta)\bigr)_{\alpha,\beta\vdash n}$ and $\bigl(H^{\bullet\hskip 1.0pt,\hskip 1.0pt\leq}(\alpha,\beta)\bigr)_{\alpha,\beta\vdash n}$ are inverse one of another, i.e.,
@@ 4.4 Relation between cyclic products and generalised precursors
for any $n$-plet of pairwise distinct indices $i_{1},i_{2},\cdots i_{n}$ and no summation over repeated indices. This relation seems to have been pointed out first in [7, Theorem 2.6], see also [8, Theorems 1 and 2]. An elegant proof has been given in a recent paper of Bernard and Hruza [9]. In fact their proof extends word for word to finite $N$, substituting the precursor $K_{n}$ for the free cumulant $\kappa_{n}$:
@@ 5.3 An illustration: Horn’s problem
Figure 1: A numerical experiment with $4\times 4$ matrices $A$ and $B$ of regularly spaced spectrum on [1,1][-1,1] and a sample of 200,000 Haar-distributed matrices $U$, showing the histograms of $\delta\kappa_{4}(U)$ (left), $\delta K_{4}(U)$ (middle) and $\delta f_{4}(U)$ (right). The improvement of the latter two with respect to the former is manifest: smaller skewness, distribution closer to Gaussian, etc. The broader range of values of $\delta K_{4}$ with respect to $\delta\kappa_{4}$ is due to the prefactor $N^{4}(N^{2}+1)/((N^{2}-1)(N^{2}-4)(N^{2}-9)$ in (55).
@@ 5.4 Behaviour of more general polynomials and the orbit coproduct
The algebra $\mathcal{A}^{(N)}$ can also be seen as the algebra of symmetric polynomials in $N$ variables, by passing to eigenvalues of matrices. We note that the algebra of symmetric polynomials in an arbitrary number of variables is equipped with a well-known cocommutative coassociative graded coproduct, defined by splitting any such polynomial into a sum of products of polynomials in two subsets of the variables (see for instance [37, Example I.5.25]). Although this coproduct shares some common properties with the one $\Delta$ studied here, it also differs from it on various aspects. To start with, $\Delta$ is defined on symmetric functions of a fixed number $N$ of variables and will depend explicitely on $N$, in contrast with this other coproduct. Moreover, the latter defines a Hopf algebra structure whereas we observed above that $\Delta$ does not. We will comment on other similarities and differences of these two coproducts throughout this subsection.
@@ 6.3 Collins–Gurau–Lionni cumulants and additive functionals
where $A$ is a random matrix which is integrated along a given measure $\mu\in\mathcal{M}^{(N)}$, while $B$ is kept deterministic and will serve as a bookkeeping variable. The Collins-Gurau-Lionni (CGL) cumulants $K^{\text{CGL}}_{\alpha}[\mu]$ are then extracted from the expansion of the logarithm of this quantity along Newton polynomials of $B$ – see [17, Equation (4.22)]: [^14] Here, we have expanded $\log\overline{Z}_{B,z}[\mu]$ along the Netwon basis $\{p_{\alpha}(B)\}_{\alpha\in\mathcal{P}}$, as defined in Section 5.4 (see above equation (113)). We have used a different normalisation than [17], which simplifies some of the discussions below.
@@
The main property of the CGL cumulants is their additivity with respect to the convolution $\star$, as established in [17, Proposition 4.9]. Namely, we have
@@ Discussion.
More generally, the normalisation in equation (144) has been chosen such that the CGL cumulants $K^{\text{CGL}}_{\alpha}$ have a finite and non-zero $N\to\infty$ limit. More precisely, according to [17, Equation (3.27)], $\lim_{N\to\infty}K^{\text{CGL}}_{\alpha}[\mu]$ becomes the higher-order free cumulant [7] associated with the partition $\alpha=(\alpha_{1},\dots,\alpha_{\ell})$ (in a certain normalisation). This is consistent with the addivity property (149) of $K^{\text{CGL}}_{\alpha}$, since higher-order free cumulants are known to be additive quantities with respect to the free convolution.
@@
[…]
@@ 7.1 Finite $\boldsymbol{N}$ free convolution associated with precursors?
$P_{A}(x)=\det(x\,\mathbb{I}-A)=\prod_{i=1}^{N}(x-\alpha_{i})=\sum_{k=0}^{N}(-1)^{k}x^{n-k}e_{k}(A)\,.\vskip-2.0pt$ (161)
@@ Acknowledgments
It is a pleasure to acknowledge stimulating discussions with D. Bernard, J. Bouttier, É. Brézin, P. Di Francesco, M. Maïda, S. Tarricone and P. Zinn-Justin. We are particularly indebted to P. Biane for drawing our attention to the point discussed in Section 7.1 and to R. Speicher for pointing out the reference [6].
@@ Appendix D Strictly monotone Hurwitz numbers and counting permutations by genus
We will need some notations and preliminary results. For $r\in\{1,\dots,n-1\}$, we let $\Sigma_{n,r}$ be the set of $r$-uples $(\tau_{1},\dots,\tau_{r})$ of transpositions in $S_{n}$ such that $\tau_{i}=(a_{i}\,b_{i})$ with $a_{i}<b_{i}$ and $b_{1}<\cdots<b_{r}$. If a permutation $\sigma\in S_{n}$ can be written as a product $\sigma=\tau_{1}\cdots\tau_{r}$ with $(\tau_{1},\dots,\tau_{r})\in\Sigma_{n,r}$, then we say that $\tau_{1}\cdots\tau_{r}$ is a primitive factorisation of $\sigma$ of length $r$. The following result is standard (see for instance [47, Lemma 1.5.1]) and will play a crucial role in our proof.